cindy charges $6.60 for 5 cartons of fries and 2 sodas and $3.82 for 3 cartons of fries and 1 soda determine the cost of one soda and one carton fries.: cindy charges $6.60 for 5 cartons of fries and 2 sodas and $3.82 for 3 cartons of fries and 1 soda determine the cost of one soda and one carton fries.
Answer by orca(336)

Let x be the unit price of fries.
Let y be the unit price of soda.
As 5 cartons of fries and 2 sodas is sold for $6.60, we have
5x + 2y = 6.6 ...........(1)
As 3 cartons of fries and 1 sodas is sold for $3.82, we have
3x + y = 3.82 ...........(2)
Next solve simultaneous equations (1) and (2) for x and y.
Multiplying equation (2) by 2, we have
6x + 2y = 7.64 ..........(3)
Subtracting equation (1) from (3), we have
x = 1.04
From equation (2), we have
y = 3.82 - 3x
Substituting x = 1.04 into it, we obtain
y = 3.82 - 3*1.04 = 0.7
So
the unit price of fries is $1.04.
the unit price of soda is $0.7